OpenStudy (anonymous):

2(x – 3) = 2x one solution no solutions infinitely many solutions 3(y – 3) = 2y – 9 + y one solution no solutions infinitely many solutions

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@dhalsted @chevygirl can you help

OpenStudy (anonymous):

1st-no solution 2nd-infinite solutions..actually dats an identity

OpenStudy (anonymous):

First expand, 2x - 6 = 2x Simplify, x - 3 = x Notice how no matter what we plug in for x, the equation won't be true? There are no solutions. First expand, 3y - 9 = 2y - 9 +y Simplify this. 3y - 9 = 3y - 9 That's an identity like srijit said, notice how you can plug anything in for y and the equation is equal, so it has infinitely many solutions.

OpenStudy (anonymous):

one last question 10x – 2 – 6x = 3x – 2 + x one solution no solutions infinitely many solutions

OpenStudy (anonymous):

@srijit

OpenStudy (anonymous):

infinite solutions

OpenStudy (anonymous):

ok thx ur awesome